A girl throws a ball 30.0 m/s [30 degrees above the horizontal], the release point of the ball is 26.0m above the bottom of a gorge. If the ball just makes it to the other side of the gorge that is 10.0m above the bottom of the gorge and we ignore air friction. What is the width of the gorge?
1
Expert's answer
2020-01-20T05:18:44-0500

"h=H+v\sin{30}t-0.5gt^2"

"10=26+30\sin{30}t-0.5(9.8)t^2"

"t=3.90\ s"

"d=v\cos{30}t=30\cos{30}3.90=101\ m"